An astronaut on the moon tosses a ball upwards; when will the ball initially be 35 feet above the lunar surface if the ball's height (feet) follow:
h(t) = -2.6t^2 = 32.1t + 5.5, and t is in seconds from release?
a. 1 second
b. 2 seconds
c. 1/2 seconds
d. 13.2 seconds
I selected 1 second (a)

Answer by stanbon(75887)

An astronaut on the moon tosses a ball upwards; when will the ball initially be 35 feet above the lunar surface if the ball's height (feet) follow:
h(t) = -2.6t^2 + 32.1t + 5.5, and t is in seconds from release?
a. 1 second
b. 2 seconds
c. 1/2 seconds
d. 13.2 seconds
-------------------
EQUATION:
-2.6t^2+32.1t + 5.5 = 35
-2.6t^2+32.1t-29.5 = 0
Since the coefficients add up to zero, x=1 is a zero.
----------
Answer: a
You are correct
=====================
Cheers,
stan H.